### v1

# LLMs as Planning Modelers: A Survey for Leveraging Large Language Models to Construct Automated Planning Models

Abstract
Large Language Models (LLMs) excel in various natural language tasks but often struggle with long-horizon planning problems requiring structured reasoning. This limitation has drawn interest in integrating neuro-symbolic approaches within the Automated Planning (AP) and Natural Language Processing (NLP) communities. However, identifying optimal AP deployment frameworks can be daunting. This paper aims to provide a timely survey of the current research with an in-depth analysis, positioning LLMs as tools for extracting and refining planning models to support reliable AP planners. By systematically reviewing the current state of research, we highlight methodologies, and identify critical challenges and future directions, hoping to contribute to the joint research on NLP and Automated Planning.
1 Introduction
Large Language Models (LLMs) has marked a significant paradigm shift in AI, sparking claims regarding emergent reasoning capabilities within LLMs (Wei et al. 2022a) and their potential integration into automated planning for agents (Pallagani et al. 2023). While LLMs, due to the prowess of distributed representation and learning, excel at System I tasks, planning—an essential aspect of System II cognition (Daniel 2017)—remains to be a significant bottleneck (Bengio 2020). Furthermore, LLMs face challenges with long-term planning and reasoning, often producing unreliable plans (Valmeekam, Stechly, and Kambhampati 2024; Pallagani et al. 2023; Momennejad et al. 2023), frequently failing to account for the effects and requirements of actions as they scale (Stechly, Valmeekam, and Kambhampati 2024), and their performance degrades with self-iterative feedback (Stechly, Marquez, and Kambhampati 2023; Valmeekam, Marquez, and Kambhampati 2023; Huang et al. 2024a).
Figure 1: Distinction of LLM-planning: (a) uses LLMs for direct planning; (b) uses LLMs to generate AP specifications for existing task planning methods such as PDDL.
Acknowledging the challenges of LLMs’ direct planning capabilities, Automated Planning (AP), or AI planning, presents a promising alternative to generate robust, optimal plans through logical and computational methods. Meanwhile, LLMs excel at extracting and refining classical planning models from Natural Language (NL), leaving the plan generation to classical planners and heuristics.
This paper is driven by the fragmented landscape in the current literature, where many surveys lack a cohesive overview of LLM integration within this field. Our focus stems from the need to address these gaps and provide a clear framework, highlighting the importance of aligning LLM capabilities with areas where they offer tangible benefits. The motivation for this approach is threefold: (i) Planning Accuracy: LLMs can help ensure that all relevant factors are considered, reducing the risk of overlooked constraints (Huang, Lipovetzky, and Cohn 2024). (ii) Adaptability: LLMs can aid systems in adapting to dynamic environments to capture real-world nuances better, reducing the need for manual edits (Dagan, Keller, and Lascarides 2023). (iii) Agnostic Modeling: LLMs enable domain-agnostic models reducing reliance on specialized expertise, a major bottleneck in AP (Gestrin, Kuhlmann, and Seipp 2024).
Our survey’s taxonomy is divided into three key areas: Model Generation, further consisting of (i) Task Generation (sec. 3.1.1), which translates initial and goal states along with object assignments; (ii) Domain Generation (sec. 3.1.2), constructing predicates and action schemas; and (iii) Hybrid Generation (sec. 3.1.3), encapsulating both task instance and domain generation. Model refinement and repair are addressed in Model Editing (sec. 3.2), while LLM-AP benchmarks are discussed in Model Benchmarks (sec. 3.3).
To our knowledge, this is the first comprehensive survey of LLM-driven AP-model construction. Our contributions can be summarized as follows:
  • A critical survey of LLM-driven Automated Planning (AP) model generation, editing, and AP-LLM benchmarks, structured within our taxonomy.
  • A summary of both shared and novel technical approaches for integrating LLMs into AI planning frameworks alongside their limitations.
  • We provide insights on key challenges and opportunities, outlining future research directions for the community. To support future work, we provide Language-to-Plan (L2P), an open-source Python library that implements landmark papers covered in this survey.
We hope this paper will contribute to and facilitate the joint research on Automated Planning and NLP.
2 Background
2.1 Automated Planning
[…]
2.2 PDDL
[…]
2.3 Large Language Models + Planning
Large Language Models (LLMs) have shown promise in generating highly structured outputs from natural language (NL) descriptions. Frameworks like StarCoder (Li et al. 2023), CodeGen (Nijkamp et al. 2023) and PLANSEARCH (Wang et al. 2024) have demonstrated LLMs’ ability to translate NL to executable code. Previous work has translated formal intermediate representations such as Temporal Logic (Chen et al. 2023b; Pan, Chou, and Berenson 2023; Mao et al. 2024; Cosler et al. 2023; Liu et al. 2023b; Luo, Xu, and Liu 2023; Fuggitti and Chakraborti 2023), Answer Set Programming (ASP) (Yang, Ishay, and Lee 2023), and other variations of the paradigm (Pan et al. 2023; Wang et al. 2023; Chen et al. 2023a). Preliminary PDDL-LLM works are found in (Miglani and Yorke-Smith 2020; Feng, Zhuo, and Kambhampati 2018; Simon and Muise 2021; Chalvatzaki et al. 2023). Currently, researchers are further exploring the nuances of varying pipelines to balance the effectiveness and limitations of LLMs in building such neuro-symbolic frameworks. Huang et al. (2024b) survey a limited amount of works to compose their high-level abstraction of LLM-augmented planning agents. Pallagani et al. (2024) go beyond the scope of traditional AP, encapsulating more broader constructs, whereas Zhao et al. (2024) provide an extensive overview of LLM-TAMP applications. Most akin to our paper, Li et al. (2024a) and Wei et al. (2025) review studies using LLMs in planning with PDDL; however, their survey mainly consists of frameworks comprised of LLMs-as-Planners, rather than LLM-as-Modelers (cf. Figure 1).
Scope of Survey: Broadly, LLMs+AP paradigms can be categorized as: (i) LLMs-as-Heuristics, where they enhance search efficiency by providing heuristic guidance (Silver et al. 2022; Tuisov, Vernik, and Shleyfman 2025; Sel, Jia, and Jin 2025); (ii) LLMs-as-Planners, either directly involved with the decision of action sequences (Zhang et al. 2024c) or proposing plans that are later refined through post-hoc implementations (Gundawar et al. 2024; Arora and Kambhampati 2023; Pallagani et al. 2022; Burns, Hughes, and Sycara 2024); and (iii) LLMs-as-Modelers, where LLMs assist in constructing planners (Katz et al. 2024; Cao et al. 2024) or planning models—defining planning domains, extracting task representations, and refining problem formulations. Our paper surveys approximately 80 papers and technical reports of the latter, focusing on key research questions and potential directions for future work in the context of LLMs and planning models.
Figure 2: Taxonomy of research in LLM Model Construction
3 Large Language Models for Constructing Planning Models
We consider the research on leveraging powerful LLMs to assist in constructing planning models to be of critical importance and a more promising approach than relying on LLMs to perform planning directly (Vyas et al. 2025). Verifiable planning modules remain the backbone of planning, ensuring reliability, robustness, and explainability—core requirements for many planning tasks. Note that deploying LLMs themselves in an end-to-end manner to perform planning still falls short of providing soundness guarantees (Valmeekam et al. 2024) and inherently struggle with correctness and completeness (Katz et al. 2024). We organize the existing works into a taxonomy comprising three key areas: Model Generation, Model Editing, and Model Benchmarks, as shown in Figure 2. These tasks need joint efforts from the NLP and automated planning community.
3.1 Model Generation
This first group of works are centered around Model Generation—extracting components or generating complete planning models from the user or environmental natural language input. This is further divided into three aspects: Task Modeling (sec. 3.1.1), which translates objectives into initial conditions and goal states; Domain Modeling (sec. 3.1.2), which defines the foundational components like entities, actions, and relationships in the system; and Hybrid Modeling (sec. 3.1.3), which integrates both aspects to create a complete model, enabling end-to-end planning.
To facilitate further discussion on LLM-driven model construction, we highlight two key research questions:
  1. RQ1
    How can LLMs accurately align with human goals, ensuring these models correctly represent desired specifications and objectives?
  2. RQ2
    To what extent and granularity of detail can human-level instructions be effectively translated into accurate model definitions?
3.1.1 Task Modeling
For goal-only specification, Collins et al. (2022) and Grover and Mohan (2024) utilize few-shot prompting whereas Faithful CoT (Lyu et al. 2023) puts heavy emphasis on an interleaving technique of chain-of-thought (CoT) prompting (Wei et al. 2022b). Xie et al. (2023) assess the effectiveness of LLMs in translating tasks with varying levels of ambiguity in both NL and other languages such as Python. Text2Motion (Lin et al. 2023) leverage LLMs to generate goal predictions that are checked against a learned library of Q-function skills—based on feasibility—that support geometric feasibility in robotic tasks. Kwon, Kim, and Kim (2024) decomposes long-term tasks into sub-goals using LLMs, then executing task planning for each sub-goal with either symbolic methods or MCTS-based LLM planners. Safe Planner (Li et al. 2024b) uses an LLM to convert NL instructions into PDDL goals, enabling a closed-loop VLM-planner to operate based on real-time environmental observations. Fine-Morris et al. (2024) propose a framework that decomposes NL goals into predicate-based specifications, structured as Python dictionaries, which are then formatted into HDDL decomposition methods.
Branching to multi-agent goal collaboration, DaTAPlan (Singh et al. 2024) employs an LLM to predict high-level anticipated tasks against human actions, triggering re-planning or new task predictions when deviations occur. PlanCollabNL (Izquierdo-Badiola et al. 2024) allocates sub-goals among LLM agents which are then encoded into PDDL goals, and modifying the action costs based on LLM recommendations. TwoStep (Singh, Traum, and Thomason 2024) decomposes multi-agent planning problems into two single-agent problems with mechanisms to ensure smooth coordination. LaMMA-P (Zhang et al. 2024d) allocate sub-tasks composed of general action sequences, and generates PDDL problem descriptions for each robot’s domain.
Frameworks that handle complete PDDL task specifications can be broadly categorized into open-loop and closed-loop approaches. For the former, LLM+P (Liu et al. 2023a) uses in-context examples of problems within the same domain to generate whole problem files. TIC (Agarwal and Sreepathy 2024) achieved nearly 100% accuracy with GPT-3.5 Turbo across LLM+P planning domains by translating the task into intermediate representations, refining them, and processing them through a logical reasoner. (Lee et al. 2024) address a critical real-world application by introducing PlanAID, a system designed to generate emergency operation plans (EOPs) by leveraging Retrieval-Augmented Generation (RAG) documents to enhance the interaction between the user and the LLM. For real-world grounding, recent work converts NL instructions into structured geometric representations that bridge abstract language understanding and spatial reasoning (Ding et al. 2023b; Chen et al. 2023a; Chang et al. 2023), while other approaches integrate LLMs with Vision Language Models (VLMs) to further ground language understanding in spatial contexts (Shirai et al. 2023; Wang et al. 2025).
For closed-loop approaches, Auto-GPT+P (Birr et al. 2024) generates the initial state of the problem based on visual perception and an automated error self-correction loop for the generated PDDL goal. Guo, Kingston, and Kavraki (2024) decompose the problem into both PDDL and Python specifications, incorporating a set of constraints into an SMT-based TAMP solver via a Python API. LLM+DP (Dagan, Keller, and Lascarides 2023) holds beliefs in uncertain environments to construct possible world states, dynamically updating its internal state and re-plans. PDDLEGO (Zhang et al. 2024a; Zhang 2024) performs a recursive task decomposition into sub-goals that enable the agent to gather new observations, progressively refining the problem file until it can develop a solvable plan. Liu et al. (2024c) integrate user information into a hierarchical scene graph of the environment, enabling an LLM to predict human activities and goal states, which are then refined using predicates and domain knowledge to ground problem specifications. Merino and Sabater-Mir (2024) model NPC behavior by leveraging LLMs conditioned on “memories” that represent environmental context.
Instead of extracting PDDL models directly, Paulius et al. (2024) introduce a modularized approach that leverages LLMs to produce Object-Level Plans (OLP), which describe high-level changes to object states and uses them to bootstrap TAMP hierarchically. TRIP-PAL (de la Rosa et al. 2024) translates intermediate representations via travel points of interest (POI) and user information into dictionaries, which are then converted into PDDL planning tasks.
Summary: Some methods directly translate natural language task descriptions into PDDL (Kelly et al. 2023; Liu et al. 2023a). Others enhance goal specification by incorporating reasoning chains and few-shot examples (Lyu et al. 2023), while (Dagan, Keller, and Lascarides 2023; Zhang et al. 2024a) target dynamic, online-planning applications. However, these current approaches rely on an explicit mapping between NL and PDDL code, requiring users to specify each state predicate in detail. To address ambiguity in minimal task descriptions, future research should develop methods capable of inferring complete and robust PDDL specifications from sparse input, building on prior work that has explored this concept via external perceptual groundings (Shirai et al. 2023), RAG implementations (Lee et al. 2024), or leveraging LLM commonsense capabilities to capture underlying assumptions and constraints (Agarwal and Sreepathy 2024).
3.1.2 Domain Modeling
Various works have executed domain modeling in a one-shot manner. Kalland (2024) combines a language and an automatic speech recognition model to generate PDDL specifications. To better understand cyber-attacks in real-time, CLLaMP (Oates et al. 2024) leverages LLMs to extract PDDL action models from Common Vulnerabilities and Exposures descriptions, finding that in-context (IC) examples are superior to CoT prompting. Zhang et al. (2024b) introduce PROC2PDDL, which proposes a Zone of Proximal Development prompt design—a variant of CoT (Vygotsky 1978). To mitigate misalignments with users requirements, Huang, Lipovetzky, and Cohn (2024) use multiple LLMs to generate candidate PDDL action schemas, which are then passed through a sentence encoder to compute the semantic relatedness of code and original NL descriptions.
Guan et al. (2023) recognize the impracticality of LLMs generating fully functioning PDDL models in a single query (Kambhampati et al. 2024). Their framework LLM+DM outlines a generate-test-critique approach (Romera-Paredes et al. 2024; Trinh et al. 2024), leveraging multiple LLM calls to incrementally build key components of the domain by a dynamically generated predicate list. Similarly, Ishay and Lee (2025) introduce LLM+AL, which uses LLMs to generate action languages, but instead, in BC+ syntax (extension of Answer Set Programming). Shah (2024) presents LAMP, an extensive series of proposed algorithms that learn abstract PDDL domain models. Sinha (2024) proposes a structured prompt engineering approach to generate domain models in the Hierarchical Planning Definition Language (HPDL). Sikes et al. (2024b) explored decomposing NL information from operational models via Javascript functions into PDDL actions. BLADE (Liu et al. 2024a) bridges language-annotated human demonstrations and primitive action interfaces by tasking an LLM to define PDDL preconditions and effects conditioned on behaviors containing all possible sequences of contact primitive.
In terms of closed-loop frameworks, ADA (Action Domain Acquisition) (Wong et al. 2023) tasks LLMs with generating candidate symbolic task decompositions, extracting undefined action names, and iteratively prompting for their definitions. COWP (Ding et al. 2023a) handles unforeseen situations in open-world planning by storing the robot’s closed-world state when planning fails, triggering a “Knowledge Acquirer” module that leverages LLMs to augment action preconditions and effects. Unlike COWP, which relies on predefined error factors, LASP (Chen et al. 2024) identifies potential errors from environmental observations, using an LLM to generate error causes in NL, suggesting action preconditions. Xie et al. (2024b) use fine-tuned LLMs for precondition and effect inference from NL actions, and semantic matching to validate actions by comparing inferred preconditions with the current world states.
Oswald et al. (2024) shifts focus to assessing domain quality, addressing limitations of manual human evaluation (Hayton et al. 2020; Huang, Chen, and Zhang 2014). This study measures equivalence to the ground truth in terms of operational equivalence—whether reconstructed domains behave identically to the original by agreeing on the validity of action sequences as plans. To achieve this, the authors decompose ground truth PDDL actions into NL using an LLM, which then tasks them again to reconstruct PDDL domain models for quality assessment.
Summary: Kambhampati et al. (2024); Wong et al. (2023) use incremental methods that iteratively generate and refine models, while incorporating real-world examples have shown to enhance contextual input accuracy (Oates et al. 2024; Ding et al. 2023a). The complexity of these frameworks demonstrate that constructing domains are inherently more challenging than task specification. However, current methods generate only a single domain, which risks misalignment with human expectations by failing to capture all necessary constraints. Future research should explore methods that generate multiple candidate domains (Huang, Lipovetzky, and Cohn 2024) and incorporate iterative refinement mechanisms to better align with user requirements.
3.1.3 Hybrid Modeling
Kelly et al. (2023) extract narrative planning domains and problems from input stories using a one-shot prompt, iterating with a second prompt conditioned on the planner’s error message until a successful plan is found. Smirnov et al. (2024) utilize pre-processing steps like JSON markup generation, consistency checks, and error correction loops. Their framework also includes a “reachability analysis” pipeline to extract feedback from flawed domains or unreachable problems, alongside a dependency analysis to check predicate usage across both files. ISR-LLM (Zhou et al. 2023) does not offer any feedback mechanisms to fix PDDL specifications; however, it does introduce self-refinement during the plan generation phase by incorporating the external validator tool, VAL (Howey, Long, and Fox 2004). Sakib and Sun (2024) generate multiple high-level task plans in Knowledge Graphs (KG), refine the network by removing unreliable components, and feed the task plan to an LLM to extract the PDDL domain and problem files for low-level robot skills. Conversely, DELTA (Liu et al. 2024b) initially generates PDDL files and a scene graph, followed by pruning unnecessary details from the graph to focus on relevant items.
Huang and Zhang (2024) further support that LLMs are prone to one-shot generation errors, highlighting the need for intermediate representations before converting to PDDL. NL2Plan (Gestrin, Kuhlmann, and Seipp 2024) is the first domain-agnostic offline end-to-end NL planning system, requiring only minimal description and using pre-processing and automated common sense feedback to interface between the LLM and the user. LLM4CAP (da Silva et al. 2024) reduces manual effort, with an LLM-generated ontology being iteratively verified using an LLM to check for syntax errors, hallucinations, and missing elements. LLMFP (Hao, Zhang, and Fan 2024) translates goals, decision variables, and constraints into a JSON representation, which is then used to generate Python code for an SMT solver to produce plans without task-specific examples or external critics. NIPE (Ying et al. 2023) leverages LLMs as few-shot semantic parsers to generate conditional statements from spatial descriptions, guiding PDDL sampling and action model definition for Bayesian goal inference.
Previous works tackle open-world settings, with MOPRHeus (Ye et al. 2024) focusing on human-in-the-loop long-horizon planning, introducing an anomaly detection mechanism to identify potential execution errors and update corresponding PDDL files to reflect changes in the world model. InterPret (Han et al. 2024) uses LLMs to enable robots to learn PDDL predicates and derive action schemas through interactive language feedback from non-expert users via Python perception APIs. Like other frameworks, Mahdavi et al. (2024) uses environmental interactions for evaluation and verification, starting with the LLM defining candidate PDDL problem files and domain sets, which are then refined through iterative cycles using their novel Exploration Walk (EW) method. Athalye et al. (2024) introduces pix2pred, a framework that leverages VLMs to propose predicates and determine their truth values in demonstrations. By bridging raw sensor data with symbolic planning, it structures the invented predicates to generalize across complex, long-horizon tasks, enabling the conversion of image data into planner-compatible representations, such as those based on PDDL.
Instead of generating planning models to be used by external planners, AgentGen (Hu et al. 2024) uses LLMs to synthesize diverse environments and planning tasks (in languages such as PDDL) and their corresponding NL descriptions, targeted for LLM-based agent training. The authors found that LLMs instruction tuned with this dataset achieved significant improvement in both in-domain and out-of-domain planning tasks. As an exploration, (Hu et al. 2025) uses this dataset to run supervised fine-tuning with the LLaMA-3.1 models and found significant improvements in generating domain models–where larger models tend to benefit more from supervised fine-tuning.
Summary: Complexities arise due to the need to coordinate between the domain and the problem. Human-in-the-loop interactions are frequently employed (Kelly et al. 2023), whereas other methods incorporate pre-processing steps–involving external tools like FastDownward and VAL (Zhou et al. 2023; Smirnov et al. 2024) or custom-designed rules (Mahdavi et al. 2024; Gestrin, Kuhlmann, and Seipp 2024). Yet, these linear pipelines risk cascading errors–such as the possibility of new objects in the task, prompting new PDDL types in the domain. Future work should focus on modularity, such as enabling dynamic integration of types and predicates in later stages of generation. This would result in more adaptable and error-tolerant planning systems.
3.2 Model Editing
Model edits have gained traction as a promising application functioning more as assistive tools than fully autonomous generative solutions. Insights into the likelihood of various model outcomes can support authors in refining models with greater efficiency toward an automated approach.
Gragera and Pozanco (2023) investigate the limitations of LLMs in repairing unsolvable tasks caused by incorrect task specifications, assessing the effectiveness of prompting in both PDDL and NL. As derivative work, Caglar et al. (2024) tackle the challenge of modifying model spaces beyond classical planning by exploring the effectiveness of LLMs in generating more plausible model edits, particularly concerning unsolvability and plan executability, either as a direct replacement or in augmentation of combinatorial search (CS) methods and manual authoring. Patil (2024) conduct a comprehensive study on using LLMs, with traditional error-checking methods, to detect and correct syntactic and semantic errors in PDDL domains, demonstrating that while LLMs excel at identifying and fixing syntactic issues, their performance in resolving semantic inconsistencies is less reliable. Sikes et al. (2024a) address the issues of planning models with domains containing a set of semantically equivalent state variables, which are syntactically different–provided with different labels. Specifically, they address failures in scenarios where information about the model is coming from heterogeneous sources.
Summary: Current research shows promise in leveraging LLMs to correct syntactic errors, but addressing semantic errors remains a significant challenge (Patil 2024). To overcome these issues, future work should explore post-hoc correction strategies. For instance, researchers could develop methodologies to analyze plan outputs after generation, identifying semantic gaps or inconsistencies through automated metrics or even human-in-the-loop evaluation systems. Such feedback can iteratively refine planning models, ensuring plans are both syntactically correct and semantically faithful.
3.3 Model Benchmarks
Non-deterministic output behaviors inherent in LLMs make it challenging to assess the quality of frameworks used in planning benchmarks. This heightens the importance of robustness, especially for evaluating LLMs’ ability to extract planning models (Behnke and Bercher 2024). However, this also introduces challenges of automating evaluation with these extracted models, determining levels of ambiguity in natural language inputs and assumptions, and monitoring fine-granular modes of failure in generation.
PlanBench (Valmeekam et al. 2023) aims to systematically evaluate LLM planning capabilities with an emphasis on cost-optimal planning and plan verification. ACPBench (Kokel et al. 2024) standardizes evaluation tasks and metrics for assessing reasoning about actions, changes (transitions), and planning—across 13 domains, on 22 SOTA language models including OpenAI’s o1 reasoning model. AutoPlanBench (Stein and Koller 2023) first converts PDDL planning benchmarks into NL via LLMs, and then tasks LLMs to produce a plan through various prompting techniques. Bohnet et al. (2024) provide a scalable benchmark suite in both PDDL and NL to measure the planning capabilities of LLMs in various strategies, as well as providing a mapping method for translating PDDL benchmarks to NL and measuring the performance of the generated benchmarks. Puerta-Merino et al. (2025) propose a road map and benchmark to address the gap of LLM integration in Hierarchical Planning (HP).
To determine whether testing PDDL domains have been leaked to training data of LLMs, Mystery Blocksworld (Kambhampati et al. 2024) obfuscates the classic Blocksworld (Gupta, Efros, and Hebert 2010) planning problem by altering the named types to be syntactically nonsensical. ALFWorld and Household (Shridhar et al. 2021; Guan et al. 2023) tackles the complexities of real-world typical household environment that uses PDDL semantics to produce textual observations. TravelPlanner (Xie et al. 2024a) assesses language models’ abilities in planning through agent-based interactions in a travel-planning environment. Zheng et al. (2024) extend this work with Natural Plan, which evaluates LLMs on realistic planning and scheduling benchmarks using APIs.
Previous LLM+AP benchmarks only focus on LLM plan generation capabilities, whereas Planetarium (Zuo et al. 2024) rigorously evaluates LLM-generated PDDL task specifications, highlighting two issues: (i) LLMs can produce valid code that misaligns with the original NL description, and (ii) evaluation sets often use NL descriptions too similar to the ground truth, reducing the task’s challenge. The benchmark assesses LLMs’ ability to generate PDDL problems across varying levels of abstraction and size. On the other hand, Text2World (Hu et al. 2025) introduces an automated pipeline for domain extraction and rigorous multi-criteria metrics that address the limitations of narrow domain scope (Liu et al. 2023a) and indirect evaluation methods in end-to-end plan assessments (Guan et al. 2023). Key metrics, including executability, structural similarity, and component-wise F1 scores, are employed while exploring state-of-the-art LLMs and fine-tuning techniques for generating and evaluating PDDL models.
Summary: Assessing the quality of LLM-generated PDDL models (Zuo et al. 2024; Oswald et al. 2024) has made significant progress toward rigorous evaluation; however, the rapid leakage of training data to LLMs remains a major challenge, with (Hu et al. 2025) reporting high contamination rates in the evaluation domains from (Guan et al. 2023). “Mystifying” the domains (Kambhampati et al. 2024) does not address LLMs’ limitations in handling real-world modeling tasks. Other concerns arise that even with “mystified” domains, LLMs may still memorize code structures, rendering string obfuscation ineffective (Chen et al. 2025). Future work should explore solutions for routinely changing benchmark standards for domains, actively involving the planning community in its ongoing refinement. Khandelwal, Sheth, and Agostinelli (2024) proposes a tool for generating diverse and complex planning domains, which could serve as a foundation for such a benchmark.
4 (L2P) Library
With the proliferation of related techniques to convert NL to PDDL, we are seeing an ever-increasing set of related methods. To bring them together under a single computational umbrella, and beyond just relating the work together conceptually as we have done thus far in this survey, we created a unified platform: L2P  [^1] Code made publicly at: https://github.com/AI-Planning/l2p, which re-implements landmark papers covered in this survey. This Python library is open source and has the capability of encapsulating the generalized version of the proposed “LLM-Modulo” framework (Kambhampati et al. 2024), which emphasizes soundness guarantees through iterative plan refining via external verifiers. While L2P embodies the core principles of the LLM-Modulo framework, which advocates for LLMs autonomously generating plan candidates themselves, it shifts the focus by facilitating the creation of PDDL files using LLMs – aligning with this paper’s paradigm. This approach allows for the integration of external verifiers and feedback critics, enabling users to modify and refine the extracted models. L2P offers three major benefits:
(i) Comprehensive Tool Suite: users can easily plug in various LLMs for streamlined extraction experiments with our extensive collection of PDDL extraction and refining tools.
(ii) Modular Design: facilitates flexible PDDL generation, allowing users to explore prompting styles and create customized pipelines.
(iii) Autonomous Capability: supports a fully autonomous pipeline, reducing the need for manual authoring.
Appendix B demonstrates several usage examples of L2P. To demonstrate the flexibility of the framework, we re-implemented several of the landmark papers covered in this survey, and Appendix C provides potential works that can be recreated by L2P at the time of writing. We hope that the community contributes to the L2P framework as a repository of existing advancements in LLM model acquisition and relevant papers, ensuring that users have access to the most current research and tools under a common framework for fair comparison.
4.1 Paper Reconstructions
L2P can recreate and encompass previous frameworks for converting natural language to PDDL, serving as a comprehensive foundation that integrates past approaches. The L2P GitHub contains multiple paper reconstructions as examples of how existing methods can be implemented and compared within a unified system, demonstrating L2P’s flexibility and effectiveness in standardizing diverse NL-to-PDDL techniques. An example of Guan et al. (2023) ”action-by-action” algorithm can be found in Figure 3; predicate output can be found in Appendix B Figure 7.
1import os
2from l2p import *
3
4def run_aba_alg(model: LLM, action_model,
5 domain_desc, hierarchy, prompt, max_iter: int=2
6 ) -> tuple[list[Predicate], list[Action]]:
7actions = list(action_model.keys())
8pred_list = []
9
10for _ in range(max_iter):
11 action_list = []
12 # iterate each action spec. + new predicates
13 for _, action in enumerate(actions):
14 if len(pred_list) == 0:
15 prompt = prompt.replace(’{predicates}’,
[…]
17 else: res = ""
18 for i, p in enumerate(pred_list):
[…]
20 prompt = prompt.replace(’{predicates}’, res)
21 # extract pddl action and predicates (L2P)
22 pddl_action, new_preds, response = (
23 builder.extract_pddl_action(
24 model=model,
25 domain_desc=domain_desc,
26 prompt_template=prompt,
27 action_name=action,
28 action_desc=action_model[action][’desc’],
29 action_list=action_list,
30 predicates=pred_list,
31 types=hierarchy["hierarchy"]
32 )
33 )
34 new_preds = parse_new_predicates(response)
35 pred_list.extend(new_preds)
36 action_list.append(pddl_action)
37 pred_list = prune_predicates(pred_list,action_list)
38return pred_list, action_list
39
40if __name__ == "__main__":
41 builder = DomainBuilder() # domain build class
42 # load in assumption files
43 base_path=’paper_reconstructions/llm+dm/prompts/’
44 action_model=load_file(
45 f’{base_path}action_model.json’)
46 domain_desc=load_file(
47 f’{base_path}domain_desc.txt’)
48 hier=load_file(
49 f’{base_path}hierarchy_requirements.json’)
50 prompt=load_file(f’{base_path}pddl_prompt.txt’)
51 # initialise LLM engine (OpenAI in this case)
52 api_key = os.environ.get(’OPENAI_API_KEY’)
53 llm = OPENAI(model="gpt-4o-mini", api_key=api_key)
54
55 # run "action-by-action" algorithm
56 pred, action = run_aba_alg(model=llm, action_model,
57 domain_desc, hier, prompt)
Figure 3: Shortened L2P reconstruction of “action-by-action algorithm” (Guan et al. 2023). This iterates through each action description and generates its PDDL specifications while maintaining a dynamically generating predicate list.
5 Discussion
Revisiting RQ1: While frameworks can generate parsable and solvable PDDL files, it remains uncertain if these specifications align with human goals. Simple domains like Blocksworld are easier to verify. Still, scaling complex domains require users to understand how LLMs generate these specifications, emphasizing the need for explainable planning for robust, transparent, and correctable outputs (Zuo et al. 2024). Corrective feedback loops have demonstrated significant improvements in resolving failing actions preconditions (Raman et al. 2024), or re-planning in case of unexpected failures during the execution of the plan (Joublin et al. 2023; Raman et al. 2022). Ensuring alignment with user goals involves breaking down PDDL model construction into pre-processing steps with human-in-the-loop feedback, as seen in frameworks like (Kelly et al. 2023). Very reminiscent of the “critics” process in the LLM-Modulo framework (Kambhampati et al. 2024), setting up a sort of external verifier checklist and using LLMs to provide feedback, is demonstrated by Gestrin, Kuhlmann, and Seipp (2024) and Smirnov et al. (2024). A strong idea is analyzing the semantic correctness of plans generated and using that as feedback to refine the LLM-generated PDDL specifications (Sakib and Sun 2024). Additionally, intermediate representation (i.e. ASP, Python, JSON) methods that are easier for LLMs to process before converting to PDDL (Agarwal and Sreepathy 2024; Smirnov et al. 2024) can also enhance accuracy. However, over time, LLMs are likely to improve, with fewer syntax errors and more logical semantic outputs as they train on more PDDL knowledge.
Revisiting RQ2: LLMs have demonstrated that they are significantly sensitive to prompting raising questions about whether they are better off functioning as machine translators or generators. Liu et al. (2023a) demonstrate that highly explicit descriptions improve translation accuracy, while Gestrin, Kuhlmann, and Seipp (2024) and Smirnov et al. (2024) leverage minimal descriptions, relying on LLMs’ internal world knowledge to enrich outputs. Huang, Lipovetzky, and Cohn (2024), Liu et al. (2023a), and Guan et al. (2023) recognize that specifying a precise predicate set in natural language is crucial and addresses the common problem of evaluating across different methods. The challenge of operating with minimal to no textual guidance beyond the initial task prompt becomes significantly more complex in online planning contexts. This underscores the importance of standardizing prompt granularity for initial generation and iterative feedback (Liu et al. 2024b). Nabizada et al. (2024) provides a promising organized and standardized paradigm of automated generated PDDL descriptions that can be applied to LLMs.
5.1 Literature Scope and Additional Outlooks
Our survey examines nearly 80 works, with over 50 integrating LLMs into the workflow for generating AP specifications. These frameworks span task (30 papers), domain (15 papers), and hybrid (15 papers) model generation. Strong emphasis on task modeling stems from the fact that many systems, already operate within well-defined domains–such as robotics. However, advancing toward more sophisticated and adaptable systems–such as dialogue agents–requires a more holistic approach, where both task objectives and domain constraints are dynamically constructed. We encourage researchers to use this survey as a foundation for exploring beyond task modeling and toward integrated domain and hybrid specification approaches, enabling more robust and scalable end-to-end planning.
6 Conclusion
The issue of extracting these planning models has long been recognized as a major barrier to the widespread adoption of planning technologies, a topic that has been explored in depth within the knowledge acquisition literature in AP for decades (Vallati and Kitchin 2020; Hendler, Tate, and Drummond 1990). Even with the emergence of LLMs, this remains a persistent challenge, introducing a new suite of obstacles. In this survey, we examine nearly 80 scholarly articles that propose their frameworks and some other subsidiary works delegating model acquisition tasks to LLMs. By identifying the research distribution and gaps within these categories, we aim to provide a higher generalization from each framework’s methodologies into broader aspects for future architectures. Additionally, we hope researchers can apply these methodologies to more advanced planning languages with the support of our L2P library–that can be found here: https://github.com/AI-Planning/l2p.
Limitations
This survey has two primary limitations. First, regarding its scope, our focus is limited to PDDL construction frameworks and related papers. Techniques remain largely unexplored in this context, and LLM capabilities in planning are still in their early stages. Due to page space constraints, we provide only a brief overview of each work rather than an exhaustive technical analysis. Additionally, our study primarily draws works published in ACL, ACM, AAAI, NeurIPS, ICAPS, COLING, CoRR, ICML, ICRA, EMNLP, and arXiv, so there is a possibility that we may have missed relevant research from other venues. Secondly, our L2P library currently supports only basic PDDL extraction tools for fully observable deterministic planning, and does not yet include tools for areas such as conditional, temporal, or numeric planning. We plan to expand the library to cover a broader range of PDDL applications, aiming to further research into the challenges LLMs encounter in these areas.
Ethics Statement
This survey does not pose any ethical issues beyond those already present in the existing literature on planning model construction via LLMs. As with any sufficiently advanced technology, there is an opportunity for misuse of the proposed L2P library (e.g., extracting actionable planning models for unethical domains). However, we view this as a pervasive issue with all of the existing methods that aim to extract planning theories from natural language.
Appendix A Additional Information on PDDL
A.1 Why PDDL?
PDDL, as a declarative language, differs fundamentally from imperative languages like Python in how problems are expressed and solved. Declarative programming specifies what needs to be achieved rather than how to achieve it, leaving execution to an external planner. In contrast, imperative programming defines explicit step-by-step instructions, requiring precise control over execution flow. While LLMs struggle with logical reasoning in imperative programming due to its sequential dependencies, they can perform well with declarative representations like PDDL. Using LLMs for PDDL model construction is not traditional code generation but rather knowledge structuring—organizing states, actions, and constraints into a formalized model.
A.2 PDDL Example
Automated Planning is a specialized field within AI that can be challenging for those unfamiliar with its principles. A key tool in classical planning is the Planning Domain Definition Language (PDDL), which models planning problems and domains. We illustrate its concepts with the Blocksworld problem, where blocks must be stacked in a specific order using actions like picking up, unstacking, and placing blocks, all while respecting constraints like moving only one block at a time or not disturbing stacked blocks.
Demonstrated below is the Blocksworld PDDL domain file $\mathbb{DF}$:
1(define (domain blocksworld)
2 (:requirements :strips)
3 (:predicates (clear ?x) (on-table ?x) (arm-empty) (holding ?x) (on ?x ?y))
4 (:action pickup
5 :parameters (?ob)
6 :precondition (and (clear ?ob) (on-table ?ob) (arm-empty))
7 :effect (and (holding ?ob) (not (clear ?ob)) (not (on-table ?ob)) (not (arm-empty)))
8 ))
Predicates define relationships or properties that can be true or false, such as (on ?x ?y) for block ?x on ?y, (ontable ?x) for ?x on the table, (clear ?x) for ?x having nothing on top, and (holding ?x) for the robot holding ?x. Actions describe possible state changes. For instance, (pick-up) contains the parameter(s) block ?ob, preconditions requiring (clear ?ob) and (ontable ?ob), and effects updating the state to reflect the robot holding ?ob, which is no longer on the table and clear.
The following is the corresponding PDDL problem/task file $\mathbb{PF}$:
1(define (problem blocksworld-problem)
2 (:domain blocksworld)
3 (:objects A B C) ; Blocks
4 (:init (ontable A) (ontable B) (on C A) (clear B) (clear C)) ; Initial state
5 (:goal (and (on A B) (on B C)))) ; Goal state
Objects represent the entities involved, such as blocks A, B, and C. The initial state defines the starting arrangement, where blocks A and B are on the table, block C is on A, and both B and C are clear. The goal state specifies the desired configuration, where block A is stacked on B, and block B is stacked on C.
Given the above PDDL domain and problem, a classical planner might generate the following plan:
1Unstack C from A
2Put C on table
3Pick up A
4Stack A on B
5Pick up B
6Stack B on C
[…]
5
Appendix B (L2P) Framework
B.1 General Library Overview
L2P provides a comprehensive suite of tools for PDDL model creation and validation. The Builder classes enable users to prompt the LLM to generate essential components of a PDDL domain, such as types, predicates, and actions, along with their parameters, preconditions, and effects. It also supports task specification, including objects, initial, and goal states that correspond to the given domain. L2P features a customized Feedback Builder class that incorporates both LLM-generated feedback and human input, or a combination of the two. Additionally, the library includes a syntax validation tool that detects common PDDL syntax errors, using this feedback to improve the accuracy of the generated models—as illustrated in Figure 6, which shows how LLM feedback can be used to refine a PDDL problem specification. L2P is now publicly available for pip installation and can be found at: https://github.com/AI-Planning/l2p
B.2 L2P Usage Example
Below are example usages using our L2P library. Full documentation can be found on our website.
1 import os
2 from l2p import *
3
4 domain_builder = DomainBuilder() # initialize Domain Builder class
5
6 # REPLACE WITH OWN API KEY
7 api_key = os.environ.get(’OPENAI_API_KEY’)
8 llm = OPENAI(model="gpt-4o-mini", api_key=api_key)
9
10 # retrieve prompt information
11 base_path=’tests/usage/prompts/domain/’
12 domain_desc = load_file(f’{base_path}blocksworld_domain.txt’)
13 extract_predicates_prompt = load_file(f’{base_path}extract_predicates.txt’)
14 types = load_file(f’{base_path}types.json’)
15 action = load_file(f’{base_path}action.json’)
16
17 # extract predicates via LLM
18 predicates, llm_output = domain_builder.extract_predicates(
19 model=llm,
20 domain_desc=domain_desc,
21 prompt_template=extract_predicates_prompt,
22 types=types,
23 nl_actions={action[’action_name’]: action[’action_desc’]}
24 )
25
26 # format key info into PDDL strings
[…]
28
[…]
30
31 -------------------------------------------------------------------------------------
32
33 ### OUTPUT
34 (holding ?a - arm ?b - block) ; true if the arm ?a is holding the block ?b
35 (on_top ?b1 - block ?b2 - block) ; true if the block ?b1 is on top of the block ?b2
36 (clear ?b - block) ; true if the block ?b is clear (no block on top of it)
37 (on_table ?b - block) ; true if the block ?b is on the table
38 (empty ?a - arm) ; true if the arm ?a is empty (not holding any block)
Figure 4: L2P usage - generating simple PDDL predicates
[…]
10
1 import os
2 from l2p import *
3
4 task_builder = TaskBuilder() # initialize Task Builder class
5 api_key = os.environ.get(’OPENAI_API_KEY’)
6 llm = OPENAI(model="gpt-4o-mini", api_key=api_key)
7
8 # load in assumptions
9 problem_desc = load_file(r’tests/usage/prompts/problem/blocksworld_problem.txt’)
10 extract_task_prompt = load_file(r’tests/usage/prompts/problem/extract_task.txt’)
11 types = load_file(r’tests/usage/prompts/domain/types.json’)
12 predicates_json = load_file(r’tests/usage/prompts/domain/predicates.json’)
13 predicates: List[Predicate] = [Predicate(**item) for item in predicates_json]
14
15 # extract PDDL task specifications via LLM
16 objects, initial_states, goal_states, llm_response = task_builder.extract_task(
17 model=llm,
18 problem_desc=problem_desc,
19 prompt_template=extract_task_prompt,
20 types=types,
21 predicates=predicates
22 )
23
24 # format key info into PDDL strings
25 objects_str = task_builder.format_objects(objects)
26 initial_str = task_builder.format_initial(initial_states)
27 goal_str = task_builder.format_goal(goal_states)
28
29 # generate task file
30 pddl_problem = task_builder.generate_task(
31 domain="blocksworld",
32 problem="blocksworld_problem",
33 objects=objects_str,
34 initial=initial_str,
35 goal=goal_str)
36
[…]
38 -------------------------------------------------------------------------------------
39 ### LLM OUTPUT
40 (define
41 (problem blocksworld_problem)
42 (:domain blocksworld)
43 (:objects
44 blue_block - object
45 red_block - object
46 yellow_block - object
47 green_block - object
48 )
49 (:init
50 (on_top blue_block red_block)
51 (on_top red_block yellow_block)
52 (on_table yellow_block)
53 (on_table green_block)
54 (clear blue_block) (clear yellow_block) (clear green_block)
55 )
56 (:goal
57 (and (on_top red_block green_block))
58 )
59 )
Figure 5: L2P usage - generating simple PDDL task specification
[…]
10
1 import os
2 from l2p import *
3
4 feedback_builder = FeedbackBuilder()
5 api_key = os.environ.get(’OPENAI_API_KEY’)
6 llm = OPENAI(model="gpt-4o-mini", api_key=api_key)
7
8 problem_desc = load_file(r’tests/usage/prompts/problem/blocksworld_problem.txt’)
9 types = load_file(r’tests/usage/prompts/domain/types.json’)
10 feedback_template = load_file(r’tests/usage/prompts/problem/feedback.txt’)
11 predicates_json = load_file(r’tests/usage/prompts/domain/predicates.json’)
12 predicates: List[Predicate] = [Predicate(**item) for item in predicates_json]
13 llm_response = load_file(r’tests/usage/prompts/domain/llm_output_task.txt’)
14
15 objects, initial, goal, feedback_response = feedback_builder.task_feedback(
16 model=llm,
17 problem_desc=problem_desc,
18 feedback_template=feedback_template,
19 feedback_type="llm",
20 predicates=predicates,
21 types=types,
22 llm_response=llm_response)
23
[…]
25 -------------------------------------------------------------------------------------
26 ### LLM OUTPUT
27 FEEDBACK:
28 1. Are any necessary objects missing?
29 All necessary objects are included based on the problem description. So: No.
30 2. Are any unnecessary objects included?
31 All objects included are relevant to the problem. Hence: No.
32 3. Are any objects defined with the wrong type?
33 All objects are correctly defined as "object". Therefore: No.
34 4. Are any unnecessary or incorrect predicates declared?
35 All predicates used in the initial state are relevant and correctly applied. Thus: No.
36 5. Are any needed or expected predicates missing from the initial state?
37 The initial state is missing the predicate for the red block being clear. Since the red block is covered by
38 the blue block, it should not be clear. Therefore: Yes.
39 6. Is anything missing from the goal state?
40 The goal state accurately reflects the desired outcome of having the red block on top of the green block. So: No.
41 7. Is anything unnecessary included in the goal description?
42 The goal description is concise and only includes what is necessary. Therefore: No.
43 8. Should any predicate be used in a symmetrical manner?
44 The predicates used do not require symmetry as they are directional in nature. So: No.
45
46 My concrete suggestions are the following:
47 - Add the predicate to indicate that the red block is not clear:
48 - (clear red_block) should be removed from the initial state since the red block is covered by the blue block.
49
50 Final output should reflect this change in the initial state:
51 ‘‘‘
52 (clear blue_block): blue block is clear
53 (clear yellow_block): yellow block is clear
54 (clear green_block): green block is clear
55 ‘‘‘
56
57 Overall, the feedback is: Yes, the initial state needs to reflect that the red block is not clear.
Figure 6: L2P usage - generating LLM-feedback on task specification
## PREDICATES
{’name’: ’truck-at’,
    ’desc’: ’true if the truck ?t is currently at location ?l’,
    ’raw’: ’(truck-at ?t - truck ?l - location): true if the truck ?t is currently at location ?l’,
    ’params’: OrderedDict([(’?t’, ’truck’), (’?l’, ’location’)]),
    ’clean’: ’(truck-at ?t - truck ?l - location): true if the truck ?t is currently at location ?l’}
{’name’: ’package-at’,
    ’desc’: ’true if the package ?p is currently at location ?l’,
    ’raw’: ’(package-at ?p - package ?l - location): true if the package ?p is currently at location ?l’,
    ’params’: OrderedDict([(’?p’, ’package’), (’?l’, ’location’)]),
    ’clean’: ’(package-at ?p - package ?l - location): true if the package ?p is currently at location ?l’}
{’name’: ’truck-holding’,
    ’desc’: ’true if the truck ?t is currently holding the package ?p’,
    ’raw’: ’(truck-holding ?t - truck ?p - package): true if the truck ?t is currently holding the package ?p’,
    ’params’: OrderedDict([(’?t’, ’truck’), (’?p’, ’package’)]),
    ’clean’: ’(truck-holding ?t - truck ?p - package): true if the truck ?t is currently holding the package ?p’}
{’name’: ’truck-has-space’,
    ’desc’: ’true if the truck ?t has space to load more packages’,
    ’raw’: ’(truck-has-space ?t - truck): true if the truck ?t has space to load more packages’,
    ’params’: OrderedDict([(’?t’, ’truck’)]),
    ’clean’: ’(truck-has-space ?t - truck): true if the truck ?t has space to load more packages’}
{’name’: ’plane-at’,
    ’desc’: ’true if the airplane ?a is located at location ?l’,
    ’raw’: ’(plane-at ?a - plane ?l - location): true if the airplane ?a is located at location ?l’,
    ’params’: OrderedDict([(’?a’, ’plane’), (’?l’, ’location’)]),
    ’clean’: ’(plane-at ?a - plane ?l - location): true if the airplane ?a is located at location ?l’}
{’name’: ’plane-holding’,
    ’desc’: ’true if the airplane ?a is currently holding the package ?p’,
    ’raw’: ’(plane-holding ?a - plane ?p - package): true if the airplane ?a is currently holding package ?p’,
    ’params’: OrderedDict([(’?a’, ’plane’), (’?p’, ’package’)]),
    ’clean’: ’(plane-holding ?a - plane ?p - package): true if the airplane ?a is currently holding package ?p’}
{’name’: ’connected-locations’,
    ’desc’: ’true if location ?l1 is directly connected to location ?l2 in city ?c’,
    ’raw’: ’(connected-locations ?l1 - location ?l2 - location ?c - city): ?l1 is connected to ?l2 in city ?c’,
    ’params’: OrderedDict([(’?l1’, ’location’), (’?l2’, ’location’), (’?c’, ’city’)]),
    ’clean’: ’(connected-locations ?l1 - location ?l2 - location ?c - city): ?l1 is connected to ?l2 in city ?c’}
Figure 7: L2P formatted predicate output from LLM (action-by-action algorithm) of Logistics domain.
Appendix C Paper Overview
This section contains an overview of the core framework papers found within this survey. The table provides a structured distinction between the papers based on their methodological approaches and specific areas of focus, such as what aspect are they modeling, prompting style/weight, and feedback mechanism. Additionally, it highlights papers already covered by the L2P library, ensuring users have quick access to resources that have been implemented or analyzed in detail. These covered papers will be included in our GitHub repository for easy reference. The distinction in the table also aims to emphasize gaps in the literature, such as under explored techniques or novel applications of NL-PDDL through LLMs, paving the way for future research. To maintain relevance, we plan to keep an updated installment of all NL-PDDL works leveraging LLMs, ensuring this survey evolves alongside the field.
Framework Task Domain NHI Prompt Feedback
(Collins et al. 2022) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) None
(Xie et al. 2023) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) None
(Lyu et al. 2023) $\checkmark$ $\times$ $\checkmark$ Light (CoT) None
(Grover and Mohan 2024) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) Env.
(Lin et al. 2023) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) LLM+Env.
(Liu et al. 2023a) $\checkmark$ $\times$ $\checkmark$ Heavy (Few-shot) None
(Birr et al. 2024) $\checkmark$ $\times$ $\checkmark$ Light (CoT) External tool
(Agarwal and Sreepathy 2024) $\checkmark$ $\times$ $\times$ Medium None
(Dagan, Keller, and Lascarides 2023) $\checkmark$ $\times$ $\checkmark$ Medium LLM + Env.
(Zhang et al. 2024a) $\checkmark$ $\times$ $\checkmark$ Medium (Few-shot) LLM
(Liu et al. 2024c) $\checkmark$ $\times$ $\times$ Extended Scene Graph Env.
(Singh et al. 2024) $\checkmark$ $\times$ $\times$ Light LLM+Deviation
(Izquierdo-Badiola et al. 2024) $\checkmark$ $\times$ $\checkmark$ Light LLM
(Singh, Traum, and Thomason 2024) $\checkmark$ $\times$ $\checkmark$ Medium (One-shot+) LLM Helper Agent
(Zhang et al. 2024d) $\checkmark$ $\times$ $\checkmark$ Light FastDownward+External Tool
(Oates et al. 2024) $\times$ $\checkmark$ $\times$ Medium (Few-shot, in-context) Human
(Zhang et al. 2024b) $\times$ $\checkmark$ $\times$ Medium (Few-shot) None
(Guan et al. 2023) $\times$ $\checkmark$ $\times$ Light Human
(Huang, Lipovetzky, and Cohn 2024) $\times$ $\checkmark$ $\checkmark$ Light Sentence Encoder Filter
(Wong et al. 2023) $\times$ $\checkmark$ $\checkmark$ Medium (Few-shot) None
(Liu et al. 2024a) $\times$ $\checkmark$ $\times$ Heavy Human
(Ding et al. 2023a) $\times$ $\checkmark$ $\checkmark$ Light LLM
(Chen et al. 2024) $\times$ $\checkmark$ $\checkmark$ Medium Env.
(Kelly et al. 2023) $\checkmark$ $\checkmark$ $\checkmark$ Medium (One-shot+) External tool
(Smirnov et al. 2024) $\checkmark$ $\checkmark$ $\checkmark$ Light LLM
(Liu et al. 2024b) $\checkmark$ $\checkmark$ $\checkmark$ Medium (One-shot+) None
(Zhou et al. 2023) $\checkmark$ $\checkmark$ $\checkmark$ Heavy (Few-shot) None
(Ye et al. 2024) $\checkmark$ $\checkmark$ $\times$ Heavy Human
(Han et al. 2024) $\checkmark$ $\checkmark$ $\times$ Light Human
(Gestrin, Kuhlmann, and Seipp 2024) $\checkmark$ $\checkmark$ $\checkmark$ Heavy (Few-shot) LLM + Human
(da Silva et al. 2024) $\checkmark$ $\checkmark$ $\checkmark$ Light LLM
(Mahdavi et al. 2024) $\checkmark$ $\checkmark$ $\checkmark$ Light LLM + Env.
(Sakib and Sun 2024) $\checkmark$ $\checkmark$ $\checkmark$ Light None
(Ying et al. 2023) $\checkmark$ $\checkmark$ $\checkmark$ Medium (Few-shot) Syntax rejection filter
Table 1: Summary of frameworks in this survey. NHI = No Human Intervention. Papers reconstructed by L2P